OpenStudy (anonymous):

What is the standard form of the number in the calculator’s display? The calculator shows eight point six e plus twelve A. 8,600,000,000,000 B. 860,000,000,000 C. 0.0000000000086 D. 0.00000000000086

OpenStudy (tkhunny):

Haven't we done this, already? 8.6 E0 = 8.6 8.6 E1 = 86 8.6 E2 = 860 8.6 E3 = 8600 Can you continue? Each "E" is a power of 10.

OpenStudy (anonymous):

I still dont understand.

OpenStudy (tkhunny):

Perhaps an expanded version? 8.6 E0 = 8.6 * 10^0 = 8.6*1 = 8.6 8.6 E1 = 8.6 * 10^1 = 8.6*10 = 86 8.6 E2 = 8.6 * 10^2 = 8.6*100 = 860 8.6 E3 = 8.6 * 10^3 = 8.6*1000 = 8600 Can you continue? Each "E" is a power of 10.